# Policy Learning with a Language Bottleneck

### v2

@@ 5 SaySelect
As a proof-of-concept, we first consider SaySelect, a simple collaborative game introduced in Hu and Sadigh (2023). Here, a speaker can see a hidden set of five balls including two blue ones and must help a listener find the two blue balls by communicating with them only via numbers, see Figure 3. In principle, the speaker and listener could converge on any bijective convention between messages and balls, and this is what we find with RL agents (multiple optimal policies). Humans, on the other hand, would prefer the intuitive mapping 1 $\rightarrow$ 1, 2 $\rightarrow$ 2, etc.  that we call the human-interpretable policy. Hu and Sadigh (2023) showed that regularizing the listener’s policy with the instruction I should select the same number as my partner, successfully guided the two RL agents to the human-interpretable policy. We investigate whether PLLB could generate a similar rule and achieve a similar interpretability on its own.
@@
Task Overview. Like in Hu and Sadigh (2023), the speaker and listener are RL agents trained with Q-Learning that receive positive rewards when the listener collects blue balls, negative rewards otherwise. After training both agents for 200 episodes, we prompt llama-2-70b-chat [^4] https://huggingface.co/meta-llama/Llama-2-70b-chat-hf with the task description and a contrastive set of high- and low-reward episodes to generate a rule $\mathcal{L}$, e.g., I should choose action 4 whenever the observation is 2, 3, 4, or 5. (gen_rule, see Figure 2).
@@ 5.1 PLLB helps learn human-interpretable policies
We first compare the four methods along two dimensions: their reward and their interpretability measured as the proportional (0 to 1) similarity between the listener’s policy and the optimal human-interpretable policy (the intuitive message-ball described above). All methods but Adversarial quickly solve the task (Figure 4a), showing that corrupted rules hinder learning. Bottleneck and InstructRL both converge on the human-interpretable policy thanks to their linguistic rules while TabularQ — deprived of such rule — does not (Figure 4b). Interestingly, we note that Adversarial policies are more interpretable than TabularQ ones, suggesting that even corrupted language may provide inductive biases towards more human-like behavior. Qualitatively, we observe that generated rules converge towards describing the human-interpretable policy (e.g.,  I should follow the strategy of choosing the same action as Agent 1, see examples in Appendix Section B.1).
@@ 5.2 PLLB can learn counter-intuitive policies
A potential confound in the preceding experiment is the possibility that PLLB helps converge to a human-interpretable policy because this is the only behavior an LM can describe. To test this hypothesis, we fixed the speaker to use a counter-intuitive policy using a random mapping between balls and messages (e.g.,  1 $\rightarrow$ 3, 5 $\rightarrow$ 2, etc.). Here, the human-interpretable policy is mis-aligned, and would lead the listener to fail or learn slower. In Figure 4c we observe that while InstructRL and Adversarial, both regularized by misaligned rules (the human-interpretable one for InstructRL, another random permutation for Adversarial), are able to eventually adapt to the fixed speaker, they converge at a slower rate than both TabularQ (no rule) and Bottleneck. This effect was more pronounced at higher values of $\lambda$. Overall, Bottleneck is the only method that flexibly adapts to a fixed speaker policy while also converging on the human-like policy when the speaker is learning, without human supervision.
@@ 6 Maze
6.1 PLLB learns more generalizable policies
Figure 5a shows that, in the Standard setting, learning a valid rule (Bottleneck) does not increase learning speed of a single maze over TabularQ. However, using a corrupted rule (Adversarial) or learning from linear features (LinearQ) slows down learning. Furthermore, Bottleneck outperforms TabularQ with respect to generalization: when we switch to a new 7x7 maze sharing the same underlying color semantics, but with a different optimal action sequences. Bottleneck leverages the learned rule and generalized to the new maze more effectively than all other agents (Figure 5b). For a a fair comparison, we start with the fully converged policy for each method (requiring 100 episodes for LinearQ). While TabularQ cannot generalize because it does not perceive colors, LinearQ does generalize faster at first, but converges slower.
The generated rules improve over time to better capture the underlying structure of the maze (e.g., if I observe BLUE, then take the NORTH action; see other examples in Appendix  B.1). Across all trials, $100\%$ of the final rules mention the red $\rightarrow$ south rule and $60\%$ uncovered the more complex blue $\rightarrow$ north-then-east rule. Finally, we find that a policy that does not update Q-values (that is, purely implemented via the LM) achieves a success rate $0\%$ within the same time limit, even when using the final rules generated from PLLB. This highlights the importance of leveraging experience that cannot be expressed with language, which text-based reasoning agents cannot do.
@@ 6.3 PLLB is more interpretable and inter-operable
Figure 6: Participants provided rules from PLLB solve new mazes faster than those given either visual or no aid (control), left. They also self-reported linguistic feedback as more useful than visual (right). * marks statistical significance at level 0.05 with two-sided Mann-Whitney test after Bonferroni correction,
@@ 7.1 Task Overview
In Builder, we construct a dataset of synthetic images with a variable number of shapes (triangle, square, circle) in different colors (magenta, blue, red, green) and locations on a 2D grid. Each image is created by a sequence of discrete actions, each representing a particular combination of shape, color, and area of the grid (e.g., [ACT12 ACT2 ACT4], where ACT2 adds a magenta circle in the lower right area of the grid). In Birds, we construct a dataset of natural bird images by selecting 5 images for each of 10 bird species from the CUB-200-2011 dataset provided by Wah et al. (2011) (see Appendix Section A.3 for species list).
@@
In Builder, because target and reconstructed images are completely defined by the action sequences that generate them, we measure task success with the Levenshtein similarity between the (sorted) corresponding action sequences. However, the listener in the Birds environment outputs images using a text-to-image model that exhibit a variety of properties. We therefore introduce reward functions corresponding to three different properties for Birds: Color, Background, and Species. To evaluate a listener’s reconstructed image $x_{r}$ with respect to a particular target image $x_{t}$ and reward function, we create a contrast set $C$ by selecting the three images in CUB-200-2011 dataset that have a different value for the target reward property (e.g., color), but are most similar with respect to values for all other annotated properties. We then define reward $r$ as $r=\sum_{c\in C}d(\textsf{CLIP}(x_{r}),\textsf{CLIP}(x_{c}))-d(\textsf{CLIP}(x_{r}),\textsf{CLIP}(x_{t}))$, where CLIP refers to the CLIP image embedding (Radford et al., 2021). For both Builder and Birds tasks, reward is reported over a separate held-out validation set.
@@
We implement the listener for Builder with BART, a neural sequence-to-sequence model pre-trained on English text by Lewis et al. (2019), which we fine-tune at each episode on a training set of (description, action sequence) pairs using descriptions provided by the speaker. Meanwhile, we use the Stable Diffusion text-to-image diffusion model as the listener for Birds.
@@
We next evaluate these base descriptions using each task’s respective reward functions, and select the (image, description) examples with the 5 highest and lowest rewards. Using the prompts shown in Section A.5, we implement gen_rule using the llama-2-70b-chat and llava models for Builder and Birds, respectively. For Birds, we observe that output rules $\mathcal{L}$ are specific to the reward function: an example rule we get for the Species reward is: “Identify the bird’s species if possible, and include any distinctive characteristics that set it apart from other birds.”, while an example rule we get for the Background reward is: “Describe the setting or background of the image, such as the presence of snow, water, or other elements”.
We implement update for both tasks by simply appending the output rule $\mathcal{L}$ to the original prompt to the speaker, as shown in Figure 2, using the speaker’s output as the new description for a given target image. We repeat this cycle of eliciting rule $\mathcal{L}$ , appending $\mathcal{L}$ to the same prompt used to generate the base instructions in order to re-label our data with new instructions, and training and evaluating the listener model for a total of 5 iterations for Builder and 3 iterations for Birds. We qualitatively show this leads to improved image reconstruction from the listener in Figure 7, and next discuss our experimental results.
Figure 8: In Builder, performance of the Bottleneck listener improves upon Baseline descriptions over time with PLLB for both From Scratch and Continual Learning training settings, unlike an (Adversarial) listener that inaccurately represents rewards.
@@ 7.2 PLLB helps speakers provide more usable instructions
For both tasks, listeners following rules generated by PLLB (Bottleneck) outperform listeners trained from uninformed (Baseline) or misinformed (Adversarial) speakers (Figures 8 and 10, top row). This holds true for all three different reward functions in Birds, as well as both training settings, with Continual Learning learning enabling Bottleneck to have an even stronger improvement over Baseline image descriptions by leveraging learned task experience. Continual Learning does not significantly help the Adversarial method, emphasizing that the linguistic rules $\mathcal{L}$ in Bottleneck do capture succesfull task strategies. However, Continual Learning is not required for Bottleneck to outperform Baseline image descriptions. The success of the From Scratch training setting can be interpreted as the speaker generating abstract rules to guide its own learning, leading to improved performance and showing evidence for a cognitive use of language.
Figure 9: In Builder, study participants reconstructed target images with higher accuracy (left), and less time (right), when provided instructions that were generated using Bottleneck vs. original Baseline descriptions.
Finally, we observe that the generated rules $\mathcal{L}$ encourage the speaker to reduce vague and general language, and describe properties relevant to the corresponding underlying reward function in Birds, although this might take several cycles to fully converge (see example rules in Appendix Section B.1). Furthermore, rules $\mathcal{L}$ become more complex over time, and across different trials we observe path dependency where different trials converge to different rules (e.g., $(x,y)$ vs. $x=0.xx,y=0.yy$ for coordinates in Builder). Overall, these rules help guide the speaker and listener agents in both tasks to iteratively improve image reconstruction over time, which can also be seen qualitatively in Figure 7.
@@ 7.3 PLLB can collaborate with human listeners
Do human listeners benefit from image descriptions provided by speakers following Bottleneck rules? We conduct a human subject study for the Builder task, where we replace the listener with study participants at the end of one iteration of PLLB, and evaluate how quickly they can reconstruct target images from a held-out test set. We recruit 20 crowdworkers on Prolific to reconstruct 5 images using descriptions from a speaker following a rule $\mathcal{L}$ generated with (Bottleneck), and 5 images using the original (Baseline) descriptions. We provided participants an interface that included a drawing canvas and buttons controlling shape types and colors (Section A.2), allowing us to use our reward function of Levenshtein similarity between (sorted) user actions and the underlying action sequences for a target image.
Figure 10: For the Birds task, rules generated after 3 iterations of Language Bottleneck help automated (top) and human (bottom) speakers generate better descriptions for naturalistic image reconstruction for three different underlying reward functions (columns). Continual learning (dark teal, red in top graphs) further improves reward over listener agents learning from scratch (dark teal, red in top graphs), but not sufficiently enough to help Adversarial outperform baseline descriptions, showing that Bottleneck rules describe rewarding behavior.
Participants using Bottleneck descriptions built target images faster and more accurately than participants using control descriptions (Bonferroni-corrected Wilcoxon signed-rank test, $p<0.05$, Figure 9). 55% of them preferred Bottleneck descriptions when asked in a post-study survey, describing them as “more direct and less ambiguous”. 10% answered that both descriptions types were equally easy to follow while remaining participants preferred Baseline descriptions because they they gave more flexibility, showing a preference not captured in our reward function.
@@ 7.4 PLLB collaborates with human speakers
The descriptions from users not provided a rule are more diverse and less focused e.g., A barbed wire is an uncomfortable stop for a bird. With PLLB rules on the other hand (e.g., Describe the bird’s coloration accurately), they generate more specific descriptions: Red and black bird on barbed wire. Bright red chest, red at top of head, black wings and beak for the same target image (see other example rules in Appendix Table 4). PLLB agents can easily transmit what they learned from experience (the rule) to humans who can then benefit from that knowledge (cultural transmission, communicative use).
@@ 8 Robot Grasp Planning
Figure 11: Results in Grasp. a) Bottleneck agents learn faster than the non-linguistic RL Baseline agents, as well as Adversarial and an Oracle zero-shot prompt of a VLM without any internal policy update. b) When faced with unseen bag designs and locations, Bottleneck agents generalize and learn faster than both Baseline (which over-relies on image details) and Oracle, ( c) shows the fixed linguistic rule used). d) Examples grasps, shaded by reward, and generated rules $\mathcal{L}$ for pairs of contrasting episodes.
Can PLLB rules improve policy learning in embodied tasks with complex action spaces such as robot planning? Existing works on language model-based agents (e.g., ReAct (Yao et al., 2022) or Reflexion (Shinn et al., 2024)) and robot policies (e.g., Code as Policies) leverage pre-trained models for action generation, but cannot be applied in domains where a (non-LM-based) policy must be updated using real-world reinforcement learning signals.
@@ 8.1 Task Overview
In the Grasp task, the AI agent is a Franka Emika Panda Arm robot that needs to lift and move a paper bag filled with unknown objects off of a shelf, as shown in Appendix Figure 15. Due to the complex action space, we follow existing work by decomposing the robot’s policy into (i) perceiving the scene and identifying an appropriate grasp location, and (ii) planning a path to this location (ten Pas et al., 2017; Kalashnikov et al., 2018).
We begin with a fixed keypoint-conditioned path-planning policy which is pre-trained on a dataset of 50 trajectories with 3D-point cloud observations using Sphinx (Sundaresan et al., 2025). This model outputs a trajectory of 7-DoF end-effector poses (x/y/z/yaw/pitch/row/gripper) for the robot to execute given the keypoint.
We use PLLB to train (via RL) a grasp prediction policy that selects a grasp keypoint (a 2-D position) given an image of the environment. We define a successful grasp as one where the robot can lift the bag without it breaking, and so set our reward to be 0 for a successful grasp, -100 if the bag is not grasped at all, and use the normalized downward torque (e.g. -20Nm) applied to the robot’s shoulder joint as a dense reward for unsuccessful grasps. Initially, the agent performs random grasps. After it observes one successful grasp, we run gen_rule with the open-source molmo-7b vision–language model with contrastive images of successful and unsuccessful grasp keypoints. This gives us a linguistic rule $\mathcal{L}$ (e.g. grasp the left side of the bag) that we can use to update the policy. We regularize the initial grasping policy toward the policy $\pi_{\mathcal{L}}$ induced from rule $\mathcal{L}$ by prompting molmo-7b with $\mathcal{L}$, then instructing it to generate random keypoints consistent with the description.
We compare our Bottleneck method with an Adversarial version (swapping high and low reward grasps), a Vanilla RL baseline that only uses image observation data, and an Oracle that always guides the policy with the rule generated from zero-shot prompting molmo-7b (no constrasting examples based on a reward).

### v3

@@ 5 SaySelect
As a proof-of-concept, we first consider SaySelect, a simple collaborative game introduced in Hu and Sadigh (2023). Here, a speaker can see a hidden set of five balls including two blue ones and must help a listener find the two blue balls by communicating with them only via numbers, see Appendix Figure 8. In principle, the speaker and listener could converge on any bijective convention associating messages to balls, and this is what we find with RL agents (multiple optimal policies). Humans, on the other hand, empirically prefer the mapping 1 $\rightarrow$ 1, 2 $\rightarrow$ 2, etc.  which we call the human-interpretable policy. Whereas Hu and Sadigh (2023) showed that regularizing the listener’s policy with the instruction I should select the same number as my partner successfully guided the two RL agents to the human-interpretable policy, we ask whether PLLB could learn a similar rule and achieve a similar interpretability on its own?
Task Overview. The speaker and listener are RL agents trained with Q-Learning that receive positive rewards when the listener collects blue balls, negative rewards otherwise. After training both agents for 200 episodes, we prompt llama-2-70b-chat [^3] https://huggingface.co/meta-llama/Llama-2-70b-chat-hf with the task description and a contrastive set of high- and low-reward episodes to generate a rule $\mathcal{L}$, e.g., I should choose action 4 whenever the observation is 2, 3, 4, or 5. (gen_rule, see Figure 2).
@@ 5.1 PLLB helps learn human-interpretable policies
We first compare the four methods along two dimensions: their reward and their interpretability measured as the proportional (0 to 1) similarity between the listener’s policy and the optimal human-interpretable policy (the intuitive message-ball described above). All methods but Adversarial quickly solve the task (Figure 3a), showing that corrupted rules hinder learning. Bottleneck and InstructRL both converge on the human-interpretable policy thanks to their linguistic rules while TabularQ — deprived of such rule — does not (Figure 3b). Interestingly, Adversarial policies are more interpretable than TabularQ ones, suggesting that even corrupted language may provide inductive biases towards more human-like behavior. Qualitatively, we observe that generated rules converge towards describing the human-interpretable policy (e.g., I should follow the strategy of choosing the same action as Agent 1, see examples in Section B.8).
@@ 5.2 PLLB can learn counter-intuitive policies
A potential confound in the preceding experiment is the possibility that PLLB converges to a human-interpretable policy because this is the only behavior an LM can describe. To test this hypothesis, we fixed the speaker to use a counter-intuitive policy using a random mapping between balls and messages (e.g., 1 $\rightarrow$ 3, 5 $\rightarrow$ 2, etc.). Here, the human-interpretable policy is mis-aligned, and would lead the listener to fail or learn slower. In Figure 3, we observe that while InstructRL and Adversarial, both regularized by misaligned rules (the human-interpretable one for InstructRL, another random permutation for Adversarial), eventually adapt to the fixed speaker, they converge at a slower rate than both TabularQ (no rule) and Bottleneck. This effect was more pronounced at higher values of $\lambda$. Overall, Bottleneck is the only method that adapts to a fixed speaker policy while also converging on the human-like policy when the speaker is learning, without human supervision.
@@ 6 Maze
6.1 PLLB improves few-shot generalization
We next evaluate PLLB’s ability to improve the few-shot generalization capabilities for policies on unseen mazes with a similar underlying structure. For a a fair comparison, we start with the fully converged policy for each method (requiring 100 episodes for LinearQ). Figure 4a shows that, in the Standard setting, learning a valid rule (Bottleneck) does not increase learning speed of a single maze over TabularQ. However, using a corrupted rule (Adversarial) or learning from linear features (LinearQ) slows down learning. Furthermore, Bottleneck outperforms TabularQ with respect to few-shot generalization: when we switch to a new 7x7 maze sharing the same underlying color semantics, but with a different optimal action sequences. Bottleneck leverages the learned rule and generalizes to the new maze more effectively than all other agents (Figure 4b). While TabularQ cannot generalize because it does not perceive colors, LinearQ does generalize faster at first, but converges slower.
The generated rules improve over time to better capture the underlying structure of the maze (e.g., if I observe BLUE, then take the NORTH action; see other examples in Appendix  B.8). Across all trials, $100\%$ of the final rules mention the red $\rightarrow$ south rule and $60\%$ uncovered the more complex blue $\rightarrow$ north-then-east rule. Finally, we find that a policy that does not update Q-values (that is, purely implemented via the LM) achieves a success rate $0\%$ within the same time limit, even when using the final rules generated from PLLB. This highlights the importance of leveraging experience that cannot be expressed with language, which text-based reasoning agents cannot do.
@@ 6.3 PLLB is more interpretable and inter-operable
Figure 5: Participants given PLLB rules solve new mazes faster than those given either visual or no aid (control). They also self-reported linguistic feedback as more useful than visual. * marks statistical significance at level 0.05 with two-sided Mann-Whitney test after Bonferroni correction.
@@ 7 Collaborative Image Reconstruction
Figure 6: Builder: Bottleneck listeners improve upon Baseline descriptions over time, unlike an (Adversarial) listener that inaccurately represents rewards. Human users reconstructed target images with b. higher accuracy, and c. less time, when provided Bottleneck instruction vs. original Baseline descriptions. Birds: PLLB rules (3 iterations) help d. automated and e. human speakers generate better descriptions for naturalistic image reconstruction for three distinct reward functions.
@@ 7.1 Task Overview
In Builder, we construct a dataset of synthetic images with a variable number of shapes (triangle, square, circle) in different colors (magenta, blue, red, green) and locations on a 2D grid. Each image is created by a sequence of discrete actions, each representing a particular combination of shape, color, and location (e.g., [ACT12 ACT2 ACT4], where ACT2 adds a magenta circle in the lower right area of the grid). In Birds, we construct a dataset of natural bird images by selecting 5 images for each of 10 bird species from the CUB-200-2011 dataset provided by Wah et al. (2011) (see Appendix Section B.4 for species list).
@@
In Builder, since target and reconstructed images are fully determined by their action sequences, we measure task success via the Levenshtein similarity between the (sorted) corresponding sequences. In contrast, the Birds listener outputs text-to-image generations with varied properties, so we define three reward functions: Color, Background, and Species. To evaluate a reconstructed image $x_{r}$ against a target $x_{t}$, we form a contrast set $C$ from three CUB-200-2011 images differing in the target property (e.g., color) but most similar in others. Reward is then $r=\sum_{c\in C}d(\textsf{CLIP}(x_{r}),\textsf{CLIP}(x_{c}))-d(\textsf{CLIP}(x_{r}),\textsf{CLIP}(x_{t}))$, where CLIP denotes the CLIP image embedding (Radford et al., 2021). For both Builder and Birds, rewards are reported on a held-out validation set.
@@
We implement the listener for Builder with a neural sequence-to-sequence model pre-trained on English text (Lewis et al., 2019), which we fine-tune at each episode on a training set of (description, action sequence) pairs using descriptions provided by the speaker. Meanwhile, we use the Stable Diffusion text-to-image diffusion model as the listener for Birds.
@@
We next evaluate these base descriptions using each task’s respective reward functions, and select the (image, description) examples with the 5 highest and lowest rewards. Using the prompts shown in Section B.7, we implement gen_rule using the llama-2-70b-chat and llava models for Builder and Birds, respectively. For Birds, we observe that output rules $\mathcal{L}$ are specific to the reward function: an example rule we get for the Species reward is: “Identify the bird’s species if possible, and include any distinctive characteristics that set it apart from other birds.”, while an example rule we get for the Background reward is: “Describe the setting or background of the image, such as the presence of snow, water, or other elements”.
We implement update for both tasks by simply appending the output rule $\mathcal{L}$ to the original prompt to the speaker, as shown in Figure 2, using the speaker’s output as the new description for a given target image. We repeat this cycle of eliciting rule $\mathcal{L}$ , appending $\mathcal{L}$ to the same prompt used to generate the base instructions in order to re-label our data with new instructions, and training and evaluating the listener model for a total of 5 iterations for Builder and 3 iterations for Birds. We qualitatively show this leads to improved image reconstruction from the listener in Appendix Figure 13, and next discuss our experimental results.
@@ 7.2 PLLB helps speakers provide more usable instructions
For both tasks, listeners following rules generated by PLLB (Bottleneck) outperform listeners trained from uninformed (Baseline) or misinformed (Adversarial) speakers (Figure 6). This holds true for all three different reward functions in Birds, as well as both training settings, with Continual Learning learning enabling Bottleneck to have an even stronger improvement over Baseline image descriptions by leveraging learned task experience. Continual Learning does not significantly help the Adversarial method, emphasizing that the linguistic rules $\mathcal{L}$ in Bottleneck do capture succesfull task strategies. However, Continual Learning is not required for Bottleneck to outperform Baseline image descriptions. The success of the From Scratch training setting can be interpreted as the speaker generating abstract rules to guide its own learning, leading to improved performance and showing evidence for a cognitive use of language.
Finally, we observe that the generated rules $\mathcal{L}$ encourage the speaker to reduce vague and general language, and describe properties relevant to the corresponding underlying reward function in Birds, although this might take several cycles to fully converge (see example rules in Appendix Section B.8). Furthermore, rules $\mathcal{L}$ become more complex over time, and across different trials we observe path dependency where different trials converge to different rules (e.g., $(x,y)$ vs. $x=0.xx,y=0.yy$ for coordinates in Builder). Overall, these rules help guide the speaker and listener agents in both tasks to iteratively improve image reconstruction over time, which can also be seen qualitatively in Appendix Figure 13.
@@ 7.3 PLLB can collaborate with human listeners
Do human listeners benefit from image descriptions provided by speakers following Bottleneck rules? We conduct a human subject study for the Builder task, where we replace the listener with study participants at the end of one iteration of PLLB, and evaluate how quickly they can reconstruct target images from a held-out test set. We recruit 20 crowdworkers on Prolific to reconstruct 5 images using descriptions from a speaker following a rule $\mathcal{L}$ generated with (Bottleneck), and 5 images using the original (Baseline) descriptions. We provided participants an interface that included a drawing canvas and buttons controlling shapes and colors (Section B.3), allowing us to use Levenshtein similarity between (sorted) user actions and the underlying action sequences for a target image as reward.
Participants using Bottleneck descriptions built target images faster and more accurately than participants using control descriptions (Wilcoxon signed-rank test, $p<0.05$, Figure 6). In a post-study survey, 55% preferred Bottleneck descriptions, describing them as “more direct and less ambiguous”. 10% found both descriptions types equally easy to follow while remaining participants preferred Baseline descriptions because they gave more flexibility, leading to a more unstructured user study. However, we note that that user preference for customization does not correlate with any notion of improved performance in our original goal of image reconstruction.
@@ 7.4 PLLB collaborates with human speakers
The descriptions from users not provided a rule are more diverse and less focused e.g., A barbed wire is an uncomfortable stop for a bird. With PLLB rules on the other hand (e.g., Describe the bird’s coloration accurately), they generate more specific descriptions: Red and black bird on barbed wire. Bright red chest, red at top of head, black wings and beak for the same target image (see other example rules in Appendix Table 4). PLLB agents can easily transmit what they learned from experience (the rule) to humans.
@@ 8 Robot Grasp Planning
Figure 7: Results in Grasp. a) Bottleneck agents learn faster than the non-linguistic RL Baseline agents, as well as Adversarial and an Oracle zero-shot prompt of a VLM without any internal policy update. b) Bottleneck agents generalize faster than both Baseline (which over-relies on image details) and Oracle to unseen bag designs and locations, (c) shows the fixed linguistic rule d) Example grasps, shaded by reward.
Can PLLB rules improve policy learning in embodied tasks with complex action spaces such as robot planning? Existing works on LM-based agents (e.g., ReAct (Yao et al., 2022) or Reflexion (Shinn et al., 2024)) and robot policies (e.g., Code as Policies) leverage pre-trained models for action generation, but cannot be applied in domains where a (non-LM-based) policy must be updated using real-world RL signals.
@@ 8.1 Task Overview
In the Grasp task, our agent is a Franka Emika Panda Arm robot that needs to lift and move a paper bag filled with unknown objects off of a shelf, as shown in Appendix Figure 12. Due to the complex action space, we follow Kalashnikov et al. (2018) and decompose the robot’s policy into (i) perceiving the scene and identifying an appropriate grasp location, and (ii) planning a path to this location.
We begin with a fixed keypoint-conditioned path-planning policy which is pre-trained on a dataset of 50 trajectories with 3D-point cloud observations using Sphinx (Sundaresan et al., 2025). This model outputs a trajectory of 7-DoF end-effector poses (x/y/z/yaw/pitch/row/gripper) for the robot to execute given the keypoint. We then use PLLB to train (via RL) a grasp prediction policy that selects a grasp keypoint (a 2-D position) given an image of the environment. We define a successful grasp as one where the robot can lift the bag without it breaking, and so set our reward to be 0 for a successful grasp, -100 if the bag is not grasped at all, and use the normalized downward torque (e.g. -20Nm) applied to the robot’s shoulder joint as a dense reward for unsuccessful grasps. Initially, the agent performs random grasps. After it observes one successful grasp, we run gen_rule with the open-source molmo-7b vision–language model with contrastive images of successful and unsuccessful grasp keypoints. This gives us a linguistic rule $\mathcal{L}$ (e.g. grasp the left side of the bag) that we can use to update the policy. We regularize the initial grasping policy toward the policy $\pi_{\mathcal{L}}$ induced from rule $\mathcal{L}$ by prompting molmo-7b while conditioning on $\mathcal{L}$, then instructing it to generate random keypoints. By executing Sphinx with these keypoints, the overall policy is updated to condition on $\mathcal{L}$.
We compare our Bottleneck method with an Adversarial version (swapping high and low reward grasps and an Oracle that always guides the policy with the rule generated from zero-shot prompting molmo-7b (no constrasting examples based on a reward). We call this Zero-Shot VLM an Oracle as it represents the VLM’s privileged prior knowledge about intuitive grasping strategies, or knowledge a human expert would also possess. We also evaluate a Vanilla RL baseline where the policy only uses image observation data to predict a grasp keypoint based on task reward after execution, which helps us isolate PLLB’s incremental value for keypoint selection.